# Denoising weak lensing mass maps with diffusion model: systematic comparison with generative adversarial network

### v1

@@ 3.2 Simulation data and survey realisms
In order to evaluate the performance of denoising with GAN and DM, we generate realistic mock convergence maps as pairs of maps with and without shape noise. To this end, we employ the $\kappa$TNG mock WL data suite (Osato et al. 2021), which consists of 10,000 pseudo-independent mock weak lensing maps with the wide source redshift range ($0.03\leq z_{s}\leq 2.6$) based on the multiple plane ray-tracing approach (Hilbert et al. 2009). The large number of realisations is suitable for training data for ML-based approaches. In this work, we employ the convergence maps with the source redshift $z_{s}=1$ unless otherwise stated. The convergence map of $\kappa$TNG is square shaped with the area of $5\times 5\,\mathrm{deg}^{2}$ and pixellated with regular $1024$ grids per dimension. Though the main suite of $\kappa$TNG is generated from IllustrisTNG hydrodynamical simulations with galaxy formation physics, we use the other suite, $\kappa$TNG-Dark, simulated with IllustrisTNG-Dark dark-matter only simulations for a fair comparison with previous studies using dark-matter only simulations. Hereafter, we refer this $\kappa$TNG-Dark suite to $\kappa$TNG for simplicity.
For efficiency and stability of numerical computations, we post-process $\kappa$TNG convergence maps. First, to increase the number of maps, we divide each map into 4 parts with equal area. The size of the map after division is halved, i.e., $2.5\times 2.5\,\mathrm{deg}^{2}$ with $512$ grids per dimension. Next, to ease the computational costs, we reduce the number of grids from $512$ to $256$ by average pooling with the pooling size $2\times 2$. The final dataset consists of 40,000 maps with $2.5\times 2.5\,\mathrm{deg}^{2}$ with $256$ grids per dimension. The angular grid size is $0.59\,\mathrm{arcmin}$. Since we smooth the maps with a large smoothing scale, the pixel size is small enough to study the denoising process.
Since the original mock convergence maps are noiseless, we add the shape noise manually. The amplitude of the shape noise is determined in Eq. (8), and we adopt the intrinsic variance $\sigma_{\varepsilon}=0.35$ and the source number density $n_{\mathrm{gal}}=20\,\mathrm{arcmin}^{-2}$, which are similar to the Subaru HSC weak lensing survey (Aihara et al. 2018). Then, we assign the noise value from the Gaussian distribution with a mean of zero and a variance of $\sigma_{\mathrm{pix}}^{2}$ to each pixel independently, and the variance per pixel is given as
@@
where $\kappa_{\mathrm{true}}$ is the true $\kappa$TNG convergence map without shape noise, and $N$ is the shape noise sampled from the normal distribution with zero mean and the variance of $\sigma_{\mathrm{pix}}^{2}$. Note that the noise realisation is independent among 40,000 mass maps.
@@
where $\kappa_{\mathrm{sm}}$ is the smoothed convergence map, $W_{\mathrm{G}}(\bm{\theta})$ is the Gaussian filter and $\theta_{\mathrm{G}}$ is the smoothing scale. Hereafter, we adopt $\theta_{\mathrm{G}}=1.5\,\mathrm{arcmin}$, which is an optimal smoothing scale for the detection of massive galaxy clusters at $z_{s}\simeq 1$ (Miyazaki et al. 2018). All the relevant fields are smoothed with this filter unless otherwise stated. In summary, our goal is to reconstruct the underlying convergence field $\kappa_{\mathrm{true}}$ from the observed noisy convergence field $\kappa_{\mathrm{obs}}$.
@@ 3.3 Training and testing processes
Here, we describe the training and testing processes of GAN and DM. First, we generate 40,000 independent noise map sampled from a Gaussian distribution and add them to $\kappa$TNG maps. Then, we randomly split the 40,000 $\kappa$TNG maps into 39,000 maps for training and 1,000 maps for testing. Finally, 40,000 pairs of noiseless and noisy convergence maps are created.
In the previous work with GAN (Shirasaki et al. 2019), 40,000 noisy convergence maps are created with bootstrapping from 200 noiseless convergence maps while noise realisations are independent among 40,000 maps. Following Shirasaki et al. (2019), we generate 40,000 maps bootstrapped from 200 realisations randomly selected from the original 40,000 $\kappa$TNG maps. However, it turns out that the denoising performance is better when the maps without bootstrapping are used, i.e., all 40,000 convergence maps are independent. Therefore, in this work, we use 40,000 independent convergence maps as fiducial data set.
For training of GAN, we adopt the regularisation parameter $\lambda=100$ in Eq. (9) with a batch size of 1, which is shown to attain better performance for the generator with a U-net architecture (see, e.g., Appendix 6.2 in Isola et al. 2017). We use Adam optimiser (Kingma & Ba 2015) for training. The initial learning rate is set to 0.0002, and after 100 epochs, the learning rate decays linearly to 0 for additional 100 epochs. In total, the model is trained for 200 epochs. We train the model with a single NVIDIA A100 GPU, and it took approximately 28 hours. In contrast, testing phase, i.e., generating 1,000 denoised maps, finishes within a few minutes.
For DM, the model has several hyper-parameters, e.g., diffusion steps, sampling steps, noise schedule, and noise variance, which are specific to its model framework. These hyper-parameters are crucial for generating high-quality outputs (Nichol & Dhariwal 2021; Chen 2023). We evaluate the performance with one of hyper-parameters changed from the fiducial one and find that the diffusion steps affect the accuracy of the denoising process the most. Based on these tests, the following hyper-parameters are adopted. The model is trained for 80 epochs with a batch size of 4 and Adam optimiser. The initial learning rate is set to $5\times 10^{-5}$, and for the last 40 epochs, it was adjusted to $1\times 10^{-5}$. The timestep of diffusion process is set to 4,000, and the variance in Eq. (16) is linearly divided from $10^{-6}$ to $10^{-2}$. We use the same NVIDIA A100 GPU, and it takes approximately 45 hours for training. For the testing process, the variance is divided in the same manner as in training, with a timestep of 1,000 and a batch size of 8. Unlike GANs, the testing took approximately 3 hours with the same GPU. It should be noted that increasing the number of timesteps results in a longer testing time. For instance, setting the timestep to 2,000 requires approximately 6 hours.
@@
For denoising the WL mass maps, it is naively expected that the input map is the noisy observed map $\kappa_{\mathrm{obs}}$ and the output map is the noiseless true map $\kappa_{\mathrm{true}}$, which we refer to as direct transformation. However, the previous study (Shirasaki et al. 2019) found better performance with the indirect translation, where the output map is the noise map $N$, and the denoised map is obtained by subtracting the estimated noise map from the observed noisy map. We show the comparison of direct and indirect transformation in Appendix B and find that the direct transformation yields better results in our setting. Therefore, the direct transformation is adopted as the default in this paper.
@@ 4.1 Visual inspection
First, we visually examine the denoised convergence maps generated by GAN and DM. Figure 1 shows one realisation of the observed noisy convergence map $\kappa_{\mathrm{obs}}$, the true noiseless convergence map $\kappa_{\mathrm{true}}$, the denoised maps with GAN and DM from the test set of 1,000 realisations. Note that normalisation is not applied to the maps in Figure 1. Overall, both GAN and DM can remove noisy features at small scales and reconstruct the mass map at large scales. The bottom panels of Figure 1 display zoom-in views around the highest peak in the true noiseless convergence map. The morphology of the peak in denoised maps with GAN and DM is different, but the peak remains after denoising. However, not all peaks in noiseless maps, in particular peaks with lower heights, cannot be reconstructed in denoised maps with GAN and DM, and even spurious peaks, which do not have corresponding peaks in the noiseless map, appear in the denoised maps. This trend in peak identification has also been reported in previous studies using GAN (Shirasaki et al. 2019). The in-depth discussions on the effects of denoising on peaks are presented in Section 4.2.
@@
[…]
@@ 4.2 Peak location reconstruction
First, we discuss the origins of the peaks in the denoised maps. In this analysis, for each test map, we use one of five networks for GAN and one of five samples for DM. For each peak in the denoised map, we search for corresponding peaks which are located within $2\theta_{\mathrm{G}}=3\,\mathrm{arcmin}$ with $\mathrm{SNR}>4$ in the corresponding noiseless true map. If there is at least one peak which satisfies the criterion, the peak is classified as the peak with counterparts. Table 1 shows the number of total peaks in the denoised map, peaks with counterparts in the corresponding noiseless map and fake peaks, i.e., peaks with no counterparts in the true map. The total numbers of peaks in denoised maps with GAN and DM are quite similar at the $10\%$ level, and roughly $55\%$ of peaks have counterparts in the true maps. The rate of peaks with counterparts is slightly higher for DM than GAN.
@@
Table 1: The average numbers of total peaks, peaks with counterparts, and fake peaks per a denoised map. The numbers are measured in all 5,000 denoised maps with GAN and DM.
GAN DM
Peaks in denoised maps 20.22 18.01
Peaks with counterparts 10.95 (54.9%) 10.02 (56.1%)
Fake peaks 9.26 (45.1%) 7.99 (43.9%)
@@
Next, we focus on the peaks in the true mass maps. In this case, we divide peaks into two classes: “surviving peaks” which have the counterparts even after denoising and “missing peaks” which lose the counterparts in the denoised map. The definition of counterparts is the same as the previous case: the peaks within $2\theta_{\mathrm{G}}$ with $\mathrm{SNR}>4$ in the corresponding denoised map. Table 2 shows the numbers of surviving and missing peaks. Similarly to the previous case, the difference between GAN and DM is minor. For both models, $60\%$ of peaks in the true maps remain in the denoised maps and $40\%$ of peaks are missing in the denoised peaks for several reasons: the peak height goes below the threshold ($\mathrm{SNR}=4$) or the peak position is misplaced. The overall performance is slightly better for GAN than DM.
@@
Table 2: The average numbers of total peaks, peaks with counterparts, and fake peaks per a true map. The numbers are measured in all 1,000 true maps.
GAN DM
Peaks in true maps 17.86
Surviving peaks 10.94 (61.8%) 10.01 (56.6%)
Missing peaks 6.92 (38.2%) 7.85 (43.4%)
@@
Finally, in order to quantify the gain of denoising, we measure the performance of the peak location reconstruction without denoising. Only in this case, the denominator of the peak SNR is the standard deviation of the corresponding noiseless map instead of that of the noisy map, because the variance of the noisy map is amplified by the shape noise. Thus, we normalise noisy maps with the standard deviation of the noiseless convergence map. Table 3 shows the reconstruction results for the noisy maps. The shape noise leads to more fake peaks, and the total number of peaks is doubled compared with the case with denoising. $39.2\%$ of peaks have counterparts in the corresponding true map. Since $54.9\%$ and $56.1\%$ have counterparts for maps denoised with GAN and DM, respectively, and fewer fake peaks are generated, denoising has a practical gain in peak reconstruction.
@@
Table 3: The average numbers of total peaks, peaks with counterparts, and fake peaks per a noisy map. The numbers are measured in all 1,000 noisy maps.
Noisy
Peaks in noisy maps 43.05
Peaks with counterparts 16.19 (39.2%)
Fake peaks 26.86 (60.8%)
@@ 4.3 Pixel-level comparison
Here, we compare the denoised convergence maps with GAN and DM with the corresponding ground-truth convergence maps on a pixel basis. Figure 4 represents the relationship between the denoised and corresponding true maps. Following Shirasaki et al. (2019), the convergence maps are normalised so that the mean is zero and the variance is unity. This normalisation allows us to discern small differences between the denoised and true maps clearly. We exclude the regions near the edges within $2\theta_{\mathrm{G}}$ due to incomplete smoothing. The noisy and denoised maps are strongly correlated, and the central value is close to the correct pixel value, though there is a statistical fluctuation. Both GAN and DM reproduce the true map within $1\sigma$ for the wide range $-1\lesssim\mathrm{SNR}\lesssim 9$.
@@
We also investigate the quantitative pixel-level difference between the denoised and ground-truth maps and measure the root mean square error (RMSE) and the Pearson correlation coefficient $\rho$. Similarly to the previous analyses, the convergence maps are normalised and the region near the edges are excluded. RMSE is defined as
$\mathrm{RMSE}\equiv\sqrt{\frac{1}{N^{2}}\sum^{N}_{i,j=1}\left(\kappa_{\mathrm{denoised}}(\bm{\theta}_{ij})-\kappa_{\mathrm{true}}(\bm{\theta}_{ij})\right)^{2}},$ (30)
where $\bm{\theta}_{ij}$ is the angular position of the pixel, the subscript $i$ ($j$) denotes the pixel index along $x$-axis ($y$-axis), and $N=244$ is the number of grids per dimension after excluding the region near the edges. It should be noted that both $\kappa_{\mathrm{denoised}}$ and $\kappa_{\mathrm{true}}$ are normalised. Pearson correlation coefficient $\rho$ is computed as
@@
Table 4 shows the RMSE and Pearson coefficients of 1,000 mean or median maps of five denoising realisations and 5,000 individual maps estimated by GAN and DM. In both metrics, DM slightly outperforms GAN. Furthermore, averaging the maps from the five realisations leads to better results than individual maps.
Table 4: RMSE and Pearson correlation coefficients $\rho$ for the denoised maps with GAN and DM. These quantities are measured for five individual samples and their mean and median. As indicated by the arrows, the smaller RMSE and larger Pearson correlation coefficients correspond to a better match with the true map. The boldface denotes the best value among the individual, mean, and median maps.
RMSE $\downarrow$ Pearson coef. $\rho$uparrow
5 samples GAN DM GAN DM
Individual 0.842 0.840 0.644 0.645
Mean 0.694 0.692 0.758 0.759
Median 0.714 0.712 0.743 0.745

### v2

@@ 3.2 Simulation data and survey realisms
[…]
Since the original mock convergence maps are noiseless, we add the shape noise manually. The amplitude of the shape noise is determined in Eq. (8), and we adopt the intrinsic variance $\sigma_{\varepsilon}=0.35$ and the source number density $n_{\mathrm{gal}}=20\,\mathrm{arcmin}^{-2}$, which are similar to the Subaru HSC weak lensing survey (Aihara2018) [^3] We present denoising results with higher source number density in Appendix B. Then, we assign the noise value from the Gaussian distribution with a mean of zero and a variance of $\sigma_{\mathrm{pix}}^{2}$ to each pixel independently, and the variance per pixel is given as
@@
[…]
@@
where $\kappa_{\mathrm{sm}}$ is the smoothed convergence map, $W_{\mathrm{G}}(\bm{\theta})$ is the Gaussian filter and $\theta_{\mathrm{G}}$ is the smoothing scale. Hereafter, we adopt $\theta_{\mathrm{G}}=1.5\,\mathrm{arcmin}$, which is an optimal smoothing scale for the detection of massive galaxy clusters at $z_{s}\simeq 1$ (Miyazaki2018). All the relevant fields are smoothed with this filter unless otherwise stated. In summary, our goal is to reconstruct the underlying convergence field $\kappa_{\mathrm{true}}$ from the observed noisy convergence field $\kappa_{\mathrm{obs}}$.
@@ 3.3 Training and testing processes
[…]
In order to reduce the dynamic range of the data and suppress numerical errors, we apply the scaling to the noisy observed map $\kappa_{\mathrm{obs}}$ and the true noiseless convergence map $\kappa_{\mathrm{true}}$ for DM to match with the range of the noise amplitude (Eq. 16). The noisy and noiselss maps are rescaled as
$\kappa\to\frac{\kappa}{\kappa_{\mathrm{max}}},$ (29)
where $\kappa_{\mathrm{max}}$ is the maximum value of the entire data set and is estimated for $\kappa_{\mathrm{obs}}$ and $\kappa_{\mathrm{true}}$ separately. Note that using the data set without rescaling can also generate the denoised convergence maps at some level, but this amplitude rescaling improves the denoising performance significantly.
For training of GAN, we adopt the regularisation parameter $\lambda=100$ in Eq. (9) with a batch size of 1, which is shown to attain better performance for the generator with a U-Net architecture (see, e.g., Appendix 6.2 in Isola2017). We use Adam optimiser (Kingma2014a) for training. The initial learning rate is set to 0.0002, and after 100 epochs, the learning rate decays linearly to 0 for additional 100 epochs. In total, the model is trained for 200 epochs. We train the model with a single NVIDIA A100 GPU, and it took approximately 28 hours. In contrast, testing phase, i.e., generating 1,000 denoised maps, finishes within a few minutes.
For DM, the model has several hyper-parameters, e.g., diffusion steps, sampling steps, noise schedule, and noise variance, which are specific to its model framework. These hyper-parameters are crucial for generating high-quality outputs (Nichol2021; Chen2023). We evaluate the performance with one of hyper-parameters changed from the fiducial one and find that the diffusion steps affect the accuracy of the denoising process the most. Based on these tests, the following hyper-parameters are adopted. The model is trained for 85 epochs with a batch size of 4 and Adam optimiser. The learning rate is set to $5\times 10^{-5}$. The timestep of the diffusion process is set to 4,000, and we adopt the noise scheduling (Eq. 16) as quadratic noise schedule. Among hyper-parameters, the noise scheduling significantly affects the results at small scales. The in-depth comparison of the noise scheduling is presented in Appendix C. We use the same NVIDIA A100 GPU, and it takes approximately 45 hours for training. For the testing process, the variance is divided in the same manner as in training, with a timestep of 2,000 and a batch size of 64. Unlike GANs, the testing for 1,000 maps takes approximately 6 hours, i.e., 22 seconds per map with the same GPU. It should be noted that increasing the number of timesteps results in a longer testing time. For instance, setting the timestep to 4,000 requires approximately 12 hours.
@@
For denoising the WL mass maps, it is naively expected that the input map is the noisy observed map $\kappa_{\mathrm{obs}}$ and the output map is the noiseless true map $\kappa_{\mathrm{true}}$, which we refer to as direct transformation. However, the previous study (Shirasaki2019) found better performance with the indirect translation, where the output map is the noise map $N$, and the denoised map is obtained by subtracting the estimated noise map from the observed noisy map. We show the comparison of direct and indirect transformation in Appendix D and find that the direct transformation yields better results in our setting. Therefore, the direct transformation is adopted as the default in this paper.
@@ 4.1 Visual inspection
First, we visually examine the denoised convergence maps generated by GAN and DM. Figure 1 shows one realisation of the noisy convergence map $\kappa_{\mathrm{obs}}$, the true noiseless convergence map $\kappa_{\mathrm{true}}$, the denoised maps with GAN and DM from the test set of 1,000 realisations. Note that normalisation is not applied to the maps in Figure 1. Overall, both GAN and DM can remove noisy features at small scales and reconstruct the mass map at large scales. The bottom panels of Figure 1 display zoom-in views around the highest peak in the true noiseless convergence map. The morphology of the peak in denoised maps with GAN and DM is different, but the peak remains after denoising. However, not all peaks in noiseless maps, in particular peaks with lower heights, cannot be reconstructed in denoised maps with GAN and DM, and even spurious peaks, which do not have corresponding peaks in the noiseless map, appear in the denoised maps. This trend in peak identification has also been reported in previous studies using GAN (Shirasaki2019). The in-depth discussions on the effects of denoising on peaks are presented in Section 4.2.
@@
As described in Section 3.3, GAN and DM generate five denoised maps from a single observed map. Figure 2 presents five samples with GAN and DM for the same realisation of the convergence field in Figure 1. Though a slight difference can be seen especially around peaks, there is no significant difference in the overall structures among the five samples. In addition, Figure 3 shows the mean and median convergence maps obtained from the five denoised maps in Figure 2. These maps indicate smooth features, which suggests that taking a mean or median smears out the small-scale transients. In particular, in this specific realisation, the standard deviation of the averaged map is different by $5\%$ from the ground truth in the individual samples, but this deviation increases to about $10\text{--}20\%$ in the averaged maps.
@@ 4.2 Peak location reconstruction
First, we discuss the origins of the peaks in the denoised maps. In this analysis, for each test map, we use one of five networks for GAN and one of five samples for DM. For each peak in the denoised map, we search for corresponding peaks which are located within $2\theta_{\mathrm{G}}=3\,\mathrm{arcmin}$ with $\mathrm{SNR}>4$ in the corresponding noiseless true map. If there is at least one peak which satisfies the criterion, the peak is classified as the peak with counterparts. Table 4.2 shows the number of total peaks in the denoised map, peaks with counterparts in the corresponding noiseless map and fake peaks, i.e., peaks with no counterparts in the true map. The total numbers of peaks in denoised maps with GAN and DM are quite similar at the $10\%$ level, and roughly $55\%$ of peaks have counterparts in the true maps. The rate of peaks with counterparts is slightly higher for DM than GAN.
[…]
@@
The average numbers of total peaks, peaks with counterparts, and fake peaks per a denoised map. The numbers are measured in all 5,000 denoised maps with GAN and DM. GAN DM Peaks in denoised maps 20.22 17.65 Peaks with counterparts 10.95 (54.9%) 10.02 (55.9%) Fake peaks 9.26 (45.1%) 7.77 (44.1%)
@@
Next, we focus on the peaks in the true mass maps. In this case, we divide peaks into two classes: “surviving peaks” which have the counterparts even after denoising and “missing peaks” which lose the counterparts in the denoised map. The definition of counterparts is the same as the previous case: the peaks within $2\theta_{\mathrm{G}}$ with $\mathrm{SNR}>4$ in the corresponding denoised map. Table 4.2 shows the numbers of surviving and missing peaks. Similarly to the previous case, the difference between GAN and DM is minor. For both models, $60\%$ of peaks in the true maps remain in the denoised maps and $40\%$ of peaks are missing in the denoised peaks for several reasons: the peak height goes below the threshold ($\mathrm{SNR}=4$) or the peak position is misplaced. The overall performance is slightly better for GAN than DM.
[…]
@@
The average numbers of total peaks, peaks with counterparts, and fake peaks per a true map. The numbers are measured in all 1,000 true maps. GAN DM Peaks in true maps 17.86 Surviving peaks 10.94 (61.8%) 9.78 (55.2%) Missing peaks 6.92 (38.2%) 8.08 (44.8%)
@@
Finally, in order to quantify the gain of denoising, we measure the performance of the peak location reconstruction without denoising. Only in this case, the denominator of the peak SNR is the standard deviation of the corresponding noiseless map instead of that of the noisy map, because the variance of the noisy map is amplified by the shape noise. Thus, we normalise noisy maps with the standard deviation of the noiseless convergence map. Table 4.2 shows the reconstruction results for the noisy maps. The shape noise leads to more fake peaks, and the total number of peaks is doubled compared with the case with denoising. $39.2\%$ of peaks have counterparts in the corresponding true map. Since $54.9\%$ and $55.9\%$ have counterparts for maps denoised with GAN and DM, respectively, and fewer fake peaks are generated, denoising has a practical gain in peak reconstruction.
[…]
@@
The average numbers of total peaks, peaks with counterparts, and fake peaks per a noisy map. The numbers are measured in all 1,000 noisy maps. Noisy Peaks in noisy maps 43.05 Peaks with counterparts 16.19 (39.2%) Fake peaks 26.86 (60.8%)
@@ 4.3 Pixel-level comparison
Here, we compare the denoised convergence maps with GAN and DM with the corresponding ground-truth convergence maps on a pixel basis. Figure 4 represents the relationship between the denoised and corresponding true maps. Following Shirasaki2019, the convergence maps are normalised so that the mean is zero and the variance is unity. This normalisation allows us to discern small differences between the denoised and true maps clearly. We exclude the regions near the edges within $2\theta_{\mathrm{G}}$ due to incomplete smoothing. The noisy and denoised maps are strongly correlated, and the central value is close to the correct pixel value, though there is a statistical fluctuation. Both GAN and DM reproduce the true map within $1\sigma$ for the wide range $1\lesssim\mathrm{SNR}\lesssim 9$.
@@
We also investigate the quantitative pixel-level difference between the denoised and ground-truth maps and measure the root mean square error (RMSE) and the Pearson correlation coefficient $\rho$. RMSE is defined as
$\mathrm{RMSE}\equiv\sqrt{\frac{1}{N^{2}}\sum^{N}_{i,j=1}\left(\kappa^{\mathrm{unnorm}}_{\mathrm{denoised}}(\bm{\theta}_{ij})-\kappa^{\mathrm{unnorm}}_{\mathrm{true}}(\bm{\theta}_{ij})\right)^{2}},$ (31)
where $\bm{\theta}_{ij}$ is the angular position of the pixel, the subscript $i$ ($j$) denotes the pixel index along $x$-axis ($y$-axis), and $N=244$ is the number of grids per dimension after excluding the region near the edges. Note that only in RMSE, unnormalised raw convergence maps are used. Next, Pearson correlation coefficient $\rho$ is computed as
@@
Table 4.3 shows the RMSE and Pearson coefficients of 1,000 mean or median maps of five denoising realisations and 5,000 individual maps estimated by GAN and DM. First, we measure the RMSE and Pearson coefficients for the noisy maps without denoising, which are labelled as “No denoising” in Table 4.3. RMSE significantly improves by denoising with GAN or DM, which demonstrates the effectiveness of the denoising methods. On the other hand, the Pearson coefficient computed from individual maps stays similar even after denoising. Taking mean or median of the maps among the five realisations results in improvement for both metrics. In terms of comparison between GAN and DM, both methods show similar results in RMSE and Pearson correlation coefficient within $1\%$ level.
[…]
RMSE and Pearson correlation coefficients $\rho$ for the denoised maps with GAN and DM. These quantities are measured for five individual samples and their mean and median. The metrics computed from noisy maps before applying denoising are also shown as “No denoising”. As indicated by the arrows, the smaller RMSE and larger Pearson correlation coefficients correspond to a better match with the true map. The boldface denotes the best value among the individual, mean, and median maps. RMSE ($\times 10^{-2}$) $\downarrow$ Pearson coef. $\rho$ $\uparrow$ 5 samples GAN DM GAN DM Individual 1.12 1.11 0.644 0.638 Mean 0.87 0.86 0.758 0.757 Median 0.90 0.89 0.743 0.742 No denoising 1.47 0.67